OpenStudy (anonymous):

Calc help! A=12000(1+r/52)^520 a) find the rate of change of A with respect to r -I found the derivative to be 12000(1+r/52)^519, why isn't that correct?

OpenStudy (andrewyates):

You're missing a zero: \[ A = 12000(1+\frac{r}{52})^{520}\\ \frac{dA}{dr} = \bigg(520*12000(1+\frac{r}{52})^{519}\bigg) * \frac{d(1+\frac{r}{52})}{dr} \\ = \bigg(6240000(1+\frac{r}{52})^{519}\bigg) * \frac{1}{52} = 120,000(1+\frac{r}{52})^{519} \]

Nnesha (nnesha):

because the derivative rule is to multiply the coefficient by the exponent and then subtract one from the e exponent \[\large\rm y= ag^2\] \[y'=2a \cdot g\] and then of course for your question chain rule would work

OpenStudy (anonymous):

thank you!!

OpenStudy (andrewyates):

No problem :)